# Transverse Momentum-Dependent Heavy-Quark Fragmentation at Next-to-Leading Order

### v1

@@ 1 Introduction
Importantly, the factorization paradigm ensures that the heavy-quark TMD FFs appear as universal building blocks across predictions for a large number of processes involving final-state heavy quarks. In ref. [1], we explicitly considered their phenomenology in heavy-hadron pair production in the back-to-back limit in $e^{+}e^{-}$ collisions and in (polarized) semi-inclusive DIS at the future Electron-Ion Collider [15], cf. ref. [16] for recent dedicated projections for polarized collinear parton distribution functions (PDFs) and FFs involving heavy quarks at the EIC. However, the applicability of heavy-quark TMD FFs is much wider: Applications of immediate phenomenological interest range from $Z$$+$hadron [17] and dihadron [18] azimuthal decorrelations in $pp$ and $pA$ collisions to hadron-in-jet transverse-momentum distributions [19, 20] and energy-energy correlators in the back-to-back limit in $e^{+}e^{-}$ [21], $ep$ and $eA$ [22, 23], and $pp$ collisions [24, 25]. In all of these cases, factorization formulas involving TMD FFs have been derived, and the NLO results we present in this paper enable one to fully account for the effect of quark masses and accurately capture their highly nontrivial interplay with transverse momenta in all of these processes.
Figure -1079: The probability density $\mathrm{d}\mathcal{P}_{Q\bar{Q}\to(Qg)\bar{Q}}$ for a gluon to split off a heavy quark in the quasi-collinear limit, which coincides with the heavy-quark TMD FF $d_{1\,Q/Q}$ for finite collinear gluon momentum fraction $\bar{z}>0$ and transverse momentum $k_{T}>0$. On the rear walls of the plot we show the integrals of $d_{1\,Q/Q}$ over one of $k_{T}$ or $z$ as predicted by the NLO calculation we perform in this paper, showcasing that the TMD regularization and renormalization procedure renders them finite despite the intricate singular behavior of the integrand. The same is true for the total integral over both $z$ and $k_{T}$, which is not shown in the plot, but at the values of the scales chosen here is equal to $\frac{\alpha_{s}C_{F}}{4\pi}\bigl{(}5-\frac{\pi}{2}-\frac{2\pi^{2}}{3}\bigr{)}$. For numerical illustration, we take $\alpha_{s}(\mu=m)=0.2$.
The heavy-quark TMD FF is intimately connected with the universal gluon splitting probability off a heavy-quark in the quasi-collinear limit [26, 27], e.g. after heavy-quark pair production has taken place,
@@
where $m$ is the heavy-quark mass, $k_{T}$ is the transverse momentum of the gluon with respect to the heavy quark, and $z$ is the collinear momentum fraction retained by the heavy quark. The quasi-collinear splitting amplitude is shown in a three-dimensional plot as a function of $k_{T}$ and the momentum fraction $\bar{z}\equiv 1-z$ carried by the gluon in figure -1079. Famously, and in contrast to the behavior of light partons, the splitting amplitude is suppressed for energetic gluons with large momentum fractions $\bar{z}\gtrsim 0.2$, an effect known as the dead cone [28]. The splitting probability is strongly peaked at small $\bar{z}$ and small $k_{T}\lesssim m$. Unlike for light partons, it eventually turns over for any finite value of $\bar{z}>0$ and vanishes (near the front right edge of the plot) as $k_{T}\to 0$, yet the simultaneous limit $\bar{z},k_{T}\to 0$ diverges. A key challenge we address in this paper is how to consistently extend this complicated singular behavior of the quasi-collinear splitting probability into the limits $k_{T}\to 0,\bar{z}\to 0$ where the gluon is unresolved. The right framework for doing so is – indeed – TMD factorization [29], which provides precise field-theoretic prescriptions to isolate, regulate, and renormalize the associated divergences, and it is precisely in this sense that the heavy-quark TMD FF is the formal generalization of eq. (1.1) to $\bar{z}\to 0$ and/or $k_{T}\to 0$ in a way that is manifestly finite in four spacetime dimensions. As a demonstration, we include on the rear walls of the three-dimensional plot in figure -1079 the integral of the heavy-quark TMD FF over one of $z$ or $k_{T}\leq m$, respectively, as predicted by our calculations in this paper, showcasing that they are indeed individually finite. Similarly, the total integral over both variables $0\leq z\leq 1$ and $0\leq k_{T}\leq m$ (and thus any integral over a subset of the plane) is also finite in the usual distributional sense, and for the indicated values of the scales simply evaluates to $\frac{\alpha_{s}C_{F}}{4\pi}\bigl{(}5-\frac{\pi}{2}-\frac{2\pi^{2}}{3}\bigr{)}$. We stress that isolating this singular behavior at fixed order in terms of a universal and finite heavy-quark TMD FF is not an academic exercise, but instead is of immediate interest to extend fixed-order subtraction schemes like that of $q_{T}$ subtractions [30] and its multi-jet generalizations [31, 32, 33, 34] to quasi-collinear limits near final-state heavy quarks. Such extensions, which can potentially yield large gains in numerical efficiency, are likewise an active topic of current research, see e.g. [35, 36, 37, 38].
@@ 2.1 Heavy-quark TMD FFs
In ref. [1] we showed that up to a factor of the total probability for $Q$ to fragment into $H$, the distribution differential in $k_{T}=P_{X,T}\sim M_{H}$ and the lightcone momentum fraction $z_{H}$ of the original quark retained by $H$ is governed by a new perturbative matching coefficient, the partonic heavy-quark TMD FF $d_{1\,Q/Q}(z,k_{T},\mu,\zeta)$,
@@
where $S$ is the universal bare TMD soft function for the $\eta$ regulator, which cancels all poles of $\eta$ and the associated $\nu$ dependence within the second line and is independent of the heavy-quark mass up to secondary quark mass effects starting at two loops. By contrast, the $\overline{\text{MS}}$ renormalization factor $Z_{\mathrm{UV}}$ for quark TMDs is independent of the mass to all orders by RG consistency with the hard matching coefficient at the hard scattering energy $Q\sim p^{-}/z\gg m\sim k_{T}\sim 1/b_{T}$. While soft subtractions (known as zero bins in the SCET literature) generally need to be accounted for in both virtual and real collinear diagrams when computing the bare collinear matrix element itself, we have explicitly verified that they lead to scaleless integrals for our choice of regulator, also in the presence of the quark mass.
@@ 2.2 Nonvalence contributions
where $C_{m}(m,\mu,\zeta)$ is generated by separately matching collinear and soft modes at the scale $\mu\sim m$ onto HQET and QCD with $n_{\ell}$ flavors, respectively [43]; explicit expressions for $C_{m}$ and other perturbative ingredients to the order required for our perturbative checks are given in appendix B.
@@
where $C_{1}$ will be the target of our calculation in section 4.3. Here we have also introduced the shorthand $\rho\equiv\bar{n}\cdot v=\sqrt{\zeta}/m$ for the boost of the hadron; in ref. [1] we argued from consistency with TMD factorization that $\chi_{1,H}$ develops an anomalous dependence on this variable governed by the Collins-Soper kernel, and we will make this expectation more explicit in section 4.1.
@@ 2.4 Consistency conditions in the light-quark limit
$D_{h/i}^{(n_{\ell}+1)}(z_{h},\mu)$ $=\sum_{j}\int\!\frac{\mathrm{d}z}{z}\,\mathcal{M}_{j/i,T}(z,m,\mu)\,D_{h/j}^{(n)}\Bigl{(}\frac{z_{h}}{z},\mu\Bigr{)}\,.$ (2.23)
@@ 3.1 Bare real-emission diagrams and distributional expansion
The two contributing real-emission diagrams in Feynman gauge are given in figure -1078 (a) and (b), where $X=g$ is a single gluon. Including the mirror diagram for (a), their contributions to $d_{1\,Q/Q}^{\mathrm{bare}\,(1)}$ evaluate to
@@
This expansion is valid here because the remaining terms in eq. (3.11) are still dimensionally regulated at $\bar{z}=0$ as $k_{T}\to 0$ i.e., the remainder of the expression is a good test function for general $\epsilon$. Since the mass enters diagram (a) only through the combination $m\bar{z}$, the action of the $\delta(\bar{z})/\eta$ term on the rest of the diagram is to reduce it to the massless limit. This in particular means that the coefficient of the $1/\eta$ pole (including its exact $\epsilon$ dependence) is equal to the one for the massless TMD FF, as it must for it to cancel the pole in the $\eta$-regulated soft function, which is independent of the mass at this order. For our implementation of dimensional regularization, the soft function is given by
@@
where $f(x,\epsilon)\sim x^{-1-\epsilon}+\mathcal{O}(x^{0})$ has a dimensionally regulated singular limit $x\to 0$ on the first left-hand side. By contrast, on the right-hand side $f(x,\epsilon)$ may be expanded in $\epsilon$ since each term in the expansion is now dimensionally regulated, whereas the explicit poles are typically isolated in the Laurent series of the total integral $F(\epsilon)$. Of course, this is precisely how relations like eq. (3.14) are derived.
@@
In this form, the double pole $1/\epsilon^{2}$ is isolated at $\bar{z}=x=0$. The term proportional to $\tfrac{1}{\epsilon}\mathcal{L}_{0}(\bar{z})$ along the $x=0$ boundary ends up canceling a corresponding term from the $\epsilon$ expansion of the first term in curly braces on the second line of eq. (3.16), such that $d_{1\,Q/Q}^{\mathrm{bare}\,(a+s)}$ is finite as $\epsilon\to 0$ for $z<1$ for any $x\geq 0$. The double pole $\tfrac{1}{\epsilon}\delta(\bar{z})\delta(k_{T}^{2})$ similarly cancels.
The second diagram in figure -1078 does not require rapidity regularization. In order to expand it in the dimensional regulator, we repeat the above procedure, where we define
@@ 3.2 Virtual contributions and UV renormalization
The virtual diagram (c) from figure -1078 is common to any collinear matrix element involving massive fermions and can be evaluated in a straightforward way using integration by contours for the $\ell^{+}$ component of the loop momentum $\ell$, Feynman parametrization for the $\ell_{\perp}$ integral, and an analytic $\ell^{-}$ integral at the end. Including the mirror diagram, the result reads
@@
Eq. (3.2) is the main result of this section. Up to the terms on the first line of the right-hand side, which are predicted by the RGE and vanish for $\zeta=\mu^{2}=m^{2}$, we have cast the heavy-quark TMD FF precisely in the form of eq. (3.19), i.e., as a two-dimensional plus distribution with nontrivial functions of $z$ and $k_{T}$ living on the respective boundaries and an overall boundary contribution at $z=1$ and $k_{T}=0$. This structure was already illustrated in figure -1079, where the partial integrals over one of the variables can immediately be read off from the respective boundary terms. Of course, the total contribution in the bulk at $z<1$ and $k_{T}>0$, which can simply be read off from the content of the $[\dots]_{+,+}$ brackets accounting for the prefactor of $1/z^{2}$, is still equal to the simple result we found in eq. (3.13) and thus to the quasi-collinear splitting probability in eq. (1.1).
@@ 3.5 Nonvalence channels at $\mathcal{O}(\alpha_{s})$
Figure -1077: Diagrams contributing to the gluon-initiated partonic heavy-quark TMD FF and the mass-dependent TMD FF matching coefficients at $\mathcal{O}(\alpha_{s})$. Heavy quark propagators and field insertions are indicated in red. The mirror diagrams for (a) and (b) are understood and are included in expressions given in the text.
For completeness we also compute the partonic TMD FFs (or TMD FF matching coefficients onto collinear distributions) for all remaining partonic channels involving heavy quarks. The relevant diagrams at $\mathcal{O}(\alpha_{s})$ are given in figure -1077, where we labeled the identified final-state “hadron” by the momentum $p$. Note that for diagrams (b) and (c) compared to the otherwise identical diagrams in figure -1078, this also changes the definition of $k_{\perp}$, as the coordinate system needs to be adjusted such that the gluon now has vanishing transverse momentum. Results for antiquarks are identical at this order. For the $g\to Q$ contribution, diagram (a) evaluates to

### v3

@@ 1 Introduction
Importantly, the factorization paradigm ensures that the heavy-quark TMD FFs appear as universal building blocks across predictions for a large number of processes involving final-state heavy quarks. In ref. [1], we explicitly considered their phenomenology in heavy-hadron pair production in the back-to-back limit in $e^{+}e^{-}$ collisions and in (polarized) semi-inclusive DIS at the future Electron-Ion Collider [15], cf. ref. [16] for recent dedicated projections for polarized collinear parton distribution functions (PDFs) and FFs involving heavy quarks at the EIC. However, the applicability of heavy-quark TMD FFs is much wider: Applications of immediate phenomenological interest range from $Z$$+$hadron [17] and dihadron [18] azimuthal decorrelations in $pp$ and $pA$ collisions to hadron-in-jet transverse-momentum distributions [19, 20], transverse momentum-like event shapes (extending the calculation of mass effects in thrust-like event shapes [21, 22, 23, 24]), and energy-energy correlators in the back-to-back limit in $e^{+}e^{-}$ [25], $ep$ and $eA$ [26, 27], and $pp$ collisions [28, 29]. In all of these cases, factorization formulas involving TMD FFs have been derived, and the NLO results we present in this paper enable one to fully account for the effect of quark masses and accurately capture their highly nontrivial interplay with transverse momenta in all of these processes.
Figure -1199: The probability density $\mathrm{d}\mathcal{P}_{Q\bar{Q}\to(Qg)\bar{Q}}$ for a gluon to split off a heavy quark in the quasi-collinear limit, which coincides with the heavy-quark TMD FF $d_{1\,Q/Q}$ up to a factor of $z^{2}$ for finite collinear gluon momentum fraction $\bar{z}>0$ and transverse momentum $k_{T}>0$. On the rear walls of the plot we show the integrals of $d_{1\,Q/Q}$ over one of $k_{T}$ or $z$ as predicted by the NLO calculation we perform in this paper, showcasing that the TMD regularization and renormalization procedure renders them finite despite the intricate singular behavior of the integrand. The same is true for the total integral over both $z$ and $k_{T}<m$, which is not shown in the plot, but at the values of the scales chosen here is equal to $\frac{\alpha_{s}C_{F}}{4\pi}\bigl{(}5-\frac{\pi}{2}-\frac{2\pi^{2}}{3}\bigr{)}\approx-0.067$. For numerical illustration, we take $\alpha_{s}(\mu=m)=0.2$.
The heavy-quark TMD FF is intimately connected with the universal gluon splitting probability off a heavy quark in the quasi-collinear limit [30, 31], e.g. after heavy-quark pair production has taken place,
@@
where $m$ is the heavy-quark mass, $k_{T}$ is the transverse momentum of the gluon with respect to the heavy quark, and $z$ is the collinear momentum fraction retained by the heavy quark. The quasi-collinear splitting amplitude is shown in a three-dimensional plot as a function of $k_{T}$ and the momentum fraction $\bar{z}\equiv 1-z$ carried by the gluon in figure -1199. Famously, and in contrast to the behavior of light partons, the splitting amplitude is suppressed for energetic gluons with large momentum fractions $\bar{z}\gtrsim 0.2$, an effect known as the dead cone [32]. The splitting probability is strongly peaked at small $\bar{z}$ and small $k_{T}\lesssim m$. Unlike for light partons, it eventually turns over for any finite value of $\bar{z}>0$ and vanishes (near the front right edge of the plot) as $k_{T}\to 0$, yet the simultaneous limit $\bar{z},k_{T}\to 0$ diverges. A key challenge we address in this paper is how to consistently extend this complicated singular behavior of the quasi-collinear splitting probability into the limits $k_{T}\to 0,\bar{z}\to 0$ where the gluon is unresolved. The right framework for doing so is – indeed – TMD factorization [33], which provides precise field-theoretic prescriptions to isolate, regulate, and renormalize the associated divergences, and it is precisely in this sense that the heavy-quark TMD FF is the formal generalization of eq. (1.1) to $\bar{z}\to 0$ and/or $k_{T}\to 0$ in a way that is manifestly finite in four spacetime dimensions. As a demonstration, we include on the rear walls of the three-dimensional plot in figure -1199 the integral of the heavy-quark TMD FF over one of $z$ or $k_{T}\leq m$, respectively, as predicted by our calculations in this paper, showcasing that they are indeed individually finite. Similarly, the total integral over both variables $0\leq z\leq 1$ and $0\leq k_{T}\leq m$ (and thus any integral over a subset of the plane) is also finite in the usual distributional sense, and for the indicated values of the scales simply evaluates to $\frac{\alpha_{s}C_{F}}{4\pi}\bigl{(}5-\frac{\pi}{2}-\frac{2\pi^{2}}{3}\bigr{)}$. We stress that isolating this singular behavior at fixed order in terms of a universal and finite heavy-quark TMD FF is not an academic exercise, but instead is of immediate interest to extend fixed-order subtraction schemes like that of $q_{T}$ subtractions [34] and its multi-jet generalizations [35, 36, 37, 38] to quasi-collinear limits near final-state heavy quarks. Such extensions, which can potentially yield large gains in numerical efficiency, are likewise an active topic of current research, see e.g. [39, 40, 41, 42].
@@ 2.1 Heavy-quark TMD FFs
In ref. [1] we showed that up to a factor of the total probability $\chi_{H}$ for $Q$ to fragment into $H$, the distribution differential in $k_{T}=P_{X,T}\sim M_{H}$ and the lightcone momentum fraction $z_{H}$ of the original quark retained by $H$ is governed by a new perturbative matching coefficient, the partonic heavy-quark TMD FF $d_{1\,Q/Q}(z,k_{T},\mu,\zeta)$,
@@
where $S$ is the universal bare TMD soft function for the $\eta$ regulator, which cancels all poles of $\eta$ and the associated $\nu$ dependence, and is independent of the heavy-quark mass up to secondary quark mass effects starting at two loops. By contrast, the $\overline{\text{MS}}$ renormalization factor $Z_{\mathrm{UV}}$ for quark TMDs is independent of the mass to all orders by RG consistency with the hard matching coefficient at the hard scattering energy $Q\sim p^{-}/z\gg m\sim k_{T}\sim 1/b_{T}$. While soft subtractions (known as zero bins in the SCET literature) generally need to be accounted for in both virtual and real collinear diagrams when computing the bare collinear matrix element itself, we have explicitly verified that they lead to scaleless integrals for our choice of regulator, also in the presence of the quark mass.
@@ 2.2 Nonvalence contributions
where $C_{m}(m,\mu,\zeta)$ is generated by separately matching collinear and soft modes at the scale $\mu\sim m$ onto HQET and QCD with $n_{\ell}$ flavors, respectively [47]; explicit expressions for $C_{m}$ and other perturbative ingredients to the order required for our perturbative checks are given in appendix B. We note that eq. (2.17) may appear counterintuitive, given that the prediction of HQET in eq. (2.18) – after assuming this counting – comes out proportional to $\delta(1-z_{H})$. The proper way to understand eq. (2.17) in this context is that it may be satisfied not just point by point in $z_{H}$ far from the endpoint (where HQET predicts that the heavy-quark TMD FF is suppressed), but in particular is also satisfied when integrating over a wide bin in $z_{H}$ from $1$ down to some bin boundary $z_{H}^{\mathrm{cut}}$ with $1-z_{H}^{\mathrm{cut}}\sim 1$, or when taking Mellin moments $z_{H}^{N}$ with $N\ll m/\Lambda_{\mathrm{QCD}}$ not large. In these latter two cases, the prediction from HQET proportional to $\delta(1-z_{H})$ can be easily understood as all bin integrals and Mellin moments becoming equal and independent of $z_{H}^{\mathrm{cut}}$ and $N$ up to corrections of $\mathcal{O}(1/m)$ [1]. We point out that the assumption in eq. (2.17) in fact also underlies the “full theory” factorization for $k_{T}\sim m$ in eq. (2.1). Lowering $1-z_{H}\ll 1$ at fixed $k_{T}\sim m$ in that case requires further resummation and – in the extreme limit – a different factorization of nonperturbative physics, which we present in Section 6.
@@
where $C_{1}$ will be the target of our calculation in section 4.3. Here we have also introduced the shorthand $\rho\equiv v^{-}=\sqrt{\zeta}/m$ for the boost of the hadron; in ref. [1] we argued from consistency with TMD factorization that $\chi_{1,H}$ develops an anomalous dependence on this variable governed by the Collins-Soper kernel, and we will make this expectation more explicit in section 4.1.
@@ 2.4 Consistency conditions in the light-quark limit
$D_{h/i}^{(n_{\ell}+1)}(z_{h},\mu)$ $=\sum_{j}\int\!\frac{\mathrm{d}z}{z}\,\mathcal{M}_{j/i,T}(z,m,\mu)\,D_{h/j}^{(n_{\ell})}\Bigl{(}\frac{z_{h}}{z},\mu\Bigr{)}\,.$ (2.23)
@@ 3.1 Bare real-emission diagrams and distributional expansion
The two contributing real-emission diagrams in Feynman gauge are given in figure -1198 (a) and (b), where $X=g$ is a single gluon. Including the mirror diagram for (a), their contributions to $d_{1\,Q/Q}^{\mathrm{bare}\,(1)}$ evaluate to
@@
This expansion is valid here because the remaining terms in eq. (3.11) are still dimensionally regulated at $\bar{z}=0$ as $k_{T}\to 0$ i.e., the remainder of the expression is a good test function for finite $\epsilon\neq 0$. Since the mass enters diagram (a) only through the combination $m\bar{z}$, the action of the $\delta(\bar{z})/\eta$ term on the rest of the diagram is to reduce it to the massless limit. This in particular means that the coefficient of the $1/\eta$ pole (including its exact $\epsilon$ dependence) is equal to the one for the massless TMD FF, as it must for it to cancel the pole in the $\eta$-regulated soft function, which is independent of the mass at this order. For our implementation of dimensional regularization, the soft function is given by
@@
where $f(x,\epsilon)\sim x^{-1-\epsilon}+\mathcal{O}(x^{0})$ has a dimensionally regulated singular limit $x\to 0$ on the first left-hand side. By contrast, on the right-hand side $f(x,\epsilon)$ may be expanded in $\epsilon$ since each term in the expansion is now regulated by the plus prescription, whereas the explicit poles are typically isolated in the Laurent series of the total integral $F(\epsilon)$. Of course, this is precisely how relations like eq. (3.14) are derived.
@@
In this form, the double pole $1/\epsilon^{2}$ is isolated at $\bar{z}=x=0$. The term proportional to $\tfrac{1}{\epsilon}\mathcal{L}_{0}(\bar{z})$ along the $x=0$ boundary ends up canceling a corresponding term from the $\epsilon$ expansion of the first term in curly braces on the second line of eq. (3.16), such that $d_{1\,Q/Q}^{\mathrm{bare}\,(a+s)}$ is finite as $\epsilon\to 0$ for $z<1$ for any $x\geq 0$. The double pole $\tfrac{1}{\epsilon^{2}}\delta(\bar{z})\delta(k_{T}^{2})$ similarly cancels.
The second diagram in figure -1198 does not require rapidity regularization. In order to expand it in the dimensional regulator, we repeat the above procedure, where we define
@@ 3.2 Virtual contributions and UV renormalization
The virtual diagram (c) from figure -1198 is common to any collinear matrix element involving massive fermions and can be evaluated in a straightforward way using integration by contours for the $\ell^{+}$ component of the loop momentum $\ell$, Feynman parametrization for the $\ell_{\perp}$ integral, and an analytic $\ell^{-}$ integral at the end. Including the mirror diagram, the result reads
@@
Eq. (3.2) is the main result of this section. Up to the terms on the first line of the right-hand side, which are predicted by the RGE and vanish for $\zeta=\mu^{2}=m^{2}$, we have cast the heavy-quark TMD FF precisely in the form of eq. (3.19), i.e., as a two-dimensional plus distribution with nontrivial functions of $z$ and $k_{T}$ living on the respective boundaries and an overall boundary contribution at $z=1$ and $k_{T}=0$. This structure was already illustrated in figure -1199, where the partial integrals over one of the variables can immediately be read off from the respective boundary terms. Of course, the total contribution in the bulk at $z<1$ and $k_{T}>0$, which can simply be read off from the content of the $[\dots]_{+,+}$ brackets accounting for the prefactor of $1/z^{2}$, is still equal to the simple result we found in eq. (3.13) and thus to the quasi-collinear splitting probability in eq. (1.1).
@@ 3.5 Nonvalence channels at $\mathcal{O}(\alpha_{s})$
Figure -1197: Diagrams contributing to the gluon-initiated partonic heavy-quark TMD FF and the mass-dependent TMD FF matching coefficients at $\mathcal{O}(\alpha_{s})$. Heavy quark propagators and field insertions are indicated in red. The mirror diagrams for (a) and (b) are understood and are included in expressions given in the text.
For completeness we also compute the partonic TMD FFs (or TMD FF matching coefficients onto collinear distributions) for all remaining partonic channels involving heavy quarks. The relevant diagrams at $\mathcal{O}(\alpha_{s})$ are given in figure -1197, where we labeled the identified final-state “hadron” by the momentum $p$. Note that for diagrams (b) and (c) compared to the otherwise identical diagrams in figure -1198, this also changes the definition of $k_{\perp}$, as the coordinate system needs to be adjusted such that the gluon now has vanishing transverse momentum. Results for antiquarks are identical at this order. For the $g\to Q$ contribution, diagram (a) evaluates to